qudratic formula (4x^2)+bx-3=0 slove for b. and one solution is -5/2
Answer by mananth(16946)

(4x^2)+bx-3=0 solve for b. and one solution is -5/2
if p & q are the roots of the equation
p+q=-b/4
let p=-5/2
(-5/2)+q=-b/4
multiply by 4
-10+4q=-b...................(1)
p*q=-3/4
(-5/2)q=-3/4
q=3/10
plug q in (1)
-10+4(3/10)=b
(-100+12)/10=b
-88/10 =b
-44/5 = b